OpenStudy (trojanpoem):

Verify if this equation is true based on physics dimensions : T = 2 pi sqrt(l/g) As T = time, l = length , g = gravity

OpenStudy (michele_laino):

hint: we have: \[\sqrt {\frac{l}{g}} = \sqrt {\frac{{{\text{meters}}}}{{{\text{meters/second}}{{\text{s}}^{\text{2}}}}}} \]

OpenStudy (trojanpoem):

I've solved it and got T^-1 as the dimension. The problem is , I can't find which answer is correct a) correct b) correct according to dimensions c) wrong d) wrong according to dimensions d or c ? and why ?

OpenStudy (michele_laino):

it is correct! namely option a)

OpenStudy (trojanpoem):

why ?

OpenStudy (trojanpoem):

Ops

OpenStudy (michele_laino):

a good method is to start from the differential equation of the motion of the pendulum

OpenStudy (trojanpoem):

m/m/s^2 = sqrt(s^2)

OpenStudy (trojanpoem):

But why b is wrong ?

OpenStudy (michele_laino):

yes! it is correct, nevertheless your reasoning is not able to justify the presence of the coefficient \(2 \pi\) option b) is wrong since we have the correct formula, and the dimensional analysis can not say anything on possible numerical coefficients which may accompany the various formulas

OpenStudy (trojanpoem):

we cannot check if 2 pi is correct therefore b is correct

OpenStudy (michele_laino):

in other words, using the dimensional analysis I am able to write this: \[T \propto \sqrt {\frac{l}{g}} \]

OpenStudy (trojanpoem):

Yeah, so correct according to dimension ?

OpenStudy (michele_laino):

no, since you have this formula: \[T = 2\pi \sqrt {\frac{l}{g}} \]

OpenStudy (trojanpoem):

a or b ?

OpenStudy (michele_laino):

I think it is option a)

OpenStudy (trojanpoem):

Doesn't a mean "We are fully sure that the equation dimensions are correct, constants are correct" ?

OpenStudy (michele_laino):

this formula: \[T \propto \sqrt {\frac{l}{g}} \] is correct according to dimensions

OpenStudy (trojanpoem):

so if the constant is not written it's correct according to dimensions if it's written , just correct ?

OpenStudy (michele_laino):

yes!

OpenStudy (trojanpoem):

THANKS.